OpenStudy (anonymous):

the roots of a quadratic polynomial is here: x=1+2^(1/2), x=1-2^(1/2), What is the quadratic function in standard form.

OpenStudy (mertsj):

If the roots are : \[1+\sqrt{2} and 1-\sqrt{2}\]

OpenStudy (mertsj):

Then the factors must be: \[[x-(1+\sqrt{2})][x-(1-\sqrt{2})]

OpenStudy (mertsj):

\[[x-(1+\sqrt{2})][x-(1-\sqrt{2})]=0 \] would be the equation.

OpenStudy (mertsj):

Then you need to simplify the left side.

OpenStudy (anonymous):

how

OpenStudy (mertsj):

Use FOIL

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

i know what FOIL is

OpenStudy (anonymous):

I end up with an answer of 2

OpenStudy (mertsj):